# Localized states, topology and anomalous Hall conductivity on a 30 degrees twisted bilayer honeycomb lattice

### v3

@@ IV Acknowledgements
Appendix A Calculation of the entanglement entropy
@@
Here we briefly describe the method we use to compute entanglement entropy, which was proposed in Refs. [20, 21] and used in a number of works later on, e.g. in [23]. We consider the model of twisted bilayer at half filling given by the Hamiltonian (1) and are interested in entanglement entropy of any subsystem from the Fig. 12. First, we compute eigenvectors $\psi$ of the Hamiltonian (1) and select only the ones corresponding to its filled states. Next, we compute ’correlation matrix’, i.e. sum of their outer products
$C_{ij}=\sum\limits_{filled}\psi(\vec{r}_{i})\psi^{\dagger}(\vec{r}_{j}).$ (10)
@@
After that we select the components whose coordinates $\vec{r}_{i,j}$ are inside the subsystem whose entanglement entropy we are interested in and thus obtain ’reduced correlation matrix’ $C_{A,ij}$. The entanglement entropy can be expressed in terms of its eigenvalues $\zeta_{m}$ as
@@
Finally, to substract the leading ’area law’ terms and extract the topological contribution we compute sum of the entanglement entropies over several figures as shown on the Fig. 12.
Appendix B Kubo formula for anomalous Hall effect
@@
In this expression, we have transformed imaginary time into Matsubara frequencies representation, but left the coordinate representation. $H_{ij}$ is the lattice Hamiltonian, and $i,j$ numerate its sites. Electric field is introduced via Peierls substitution, which in turn may be approximated as
@@
$\vec{J}_{ij,\Omega}=ic^{\dagger}_{i,w+\Omega}H_{ij}(\vec{r}_{j}-\vec{r}_{i})c_{j,w}$ (14)
Note, that the current is defined here as a vector for each bond connecting the lattice sites $i$, $j$.
Now we apply Kubo formula and write that the expectation current has a form
$J_{x,ij,\Omega}=\langle J_{x,ij,\Omega}\sum\limits_{k.l}J_{y,kl,-\Omega}\rangle A_{y,kl,\Omega}.$ (15)
Now we assume that the electric field is constant in space, and the vector potential is written in the gauge $\vec{A}=\vec{E}/(i\Omega)$. The anomalous Hall conductivity is given by the antisymmetric part of the current-current correlator, namely
@@
Note that in our notations, the conductivity has just two indices $i,j$, which physically describe electric current at the bond connecting lattice sites $i,j$ as a response to constant electric field (if the electric field was not constant, we would have to include 4 components).
We evaluate the current-current correlator in a conventional way by substituting fermionic Green’s functions and summing over Matsubara frequencies. The Green’s function can be written in terms of eigenstates $\psi_{n,i}$ and energies $E_{n}$ (here $n$ numerates the states) as
$G_{ij,w}=\sum\limits_{n}\frac{\psi_{i}\psi^{\dagger}_{j}}{iw-E_{n}}.$ (17)
@@
Finally, let us look at the sum $\sum\limits_{ij}\sigma^{(Hall)}_{ij}$ over all lattice sites $i,j$. We can simplify the Eq. (II.3.3) by using explicitly the fact that $\psi$ are eigenvectors, namely by applying an identity
@@
In this way we can obtain that the total sum has the form
$\sum\limits_{i,j}\sigma^{(Hall)}_{ij}=\frac{1}{2}\sum\limits_{f,e}\sum\limits_{i,k}$ (19)
@@
After applying the completeness relation
@@
Appendix C Anomalous Hall conductivity on an infinite crystalline lattice
The goal of this section is to show that the expression for Hall conductivity (II.3.3) from the main text in the case of an infinite crystalline lattice is indeed equivalent to a well-known expression in terms of Berry curvature. Here, for clarity we separate lattice indices $i,j.$... into pairs $(i,\alpha),(j,\beta),$..., which refer to unit cells and sublattice degrees of freedom respectively.
@@ Appendix C Anomalous Hall conductivity on an infinite crystalline lattice
$\psi_{i\alpha}=\frac{1}{\sqrt{N}}e^{i\vec{k}\vec{x}_{i}}\phi_{\alpha,k}.$ (22)
Here the wavefunctions $\phi_{\alpha,k}$ are eigenfunctions of the Schrodinger equation in momentum representation
$E\phi_{\alpha,k}=\sum\limits_{\beta}H_{\alpha\beta}(k)\phi_{\beta,k}$ (23)
@@
$H_{\alpha\beta}(k)=\sum\limits_{j,\beta}e^{i\vec{k}(\vec{x}_{j}-\vec{x}_{i})}H_{i,j\atop\alpha,\beta}.$ (24)
@@
$H_{i,j\atop\alpha,\beta}=\frac{1}{N}\sum\limits_{p}H_{\alpha\beta}(p)e^{-i\vec{p}(\vec{x}_{j}-\vec{x}_{i})}.$ (25)
@@
$\sum\limits_{j}\sigma^{(Hall)}_{i,j\atop\alpha,\beta}$ $=$ $\frac{1}{2}\sum\limits_{\gamma,\delta}\frac{1}{N}\sum\limits_{\vec{k}}\frac{1}{(E_{f}-E_{E})^{2}}$
$\times\left\{\phi^{\dagger}_{e,\alpha}\frac{\partial H_{\alpha\beta}}{\partial k_{x}}\phi_{f,\beta}\cdot\phi^{\dagger}_{f,\gamma}\frac{\partial H_{\gamma\delta}(k)}{\partial k_{y}}\phi_{e,\delta}\right$.
$\qquad-\phi^{\dagger}_{f,\alpha}\frac{\partial H_{\alpha\beta}}{\partial k_{x}}\phi_{e,\beta}\cdot\phi^{\dagger}_{e,\gamma}\frac{\partial H_{\gamma\delta}(k)}{\partial k_{y}}\phi_{f,\delta}$
@@
Notice that in the left side of this equation we have sum $\sum\limits_{j}\sigma_{ij}$ of conductivities through all bonds adjacent to a site $i$ - the quantity we discuss in the main text. In addition, to obtain the physical answer we have to sum over all sublattice degrees of freedom, i.e. over the indices $\alpha,\beta$. We also note that in the Eq. (C) we still assume that the lattice is discreet, but the momentum is continuum because the lattice is infinite. Hence it is straightforward to replace the momentum summation with an integration
$\frac{1}{N}\sum\limits_{k}\to S\int\frac{d^{2}k}{(2\pi)^{2}},$ (27)
@@
$\sum\limits_{\alpha\beta}\phi^{\dagger}_{e,\alpha}\frac{\partial H_{\alpha\beta}}{\partial k_{x}}\phi_{f,\beta}=\sum\limits_{\alpha}(E_{f}-E_{e})\phi^{\dagger}_{e,\alpha}\frac{\partial\phi_{f,\alpha}}{\partial k_{x}}$

### v4

@@ IV Acknowledgements
Appendix A The limit of $30^{\circ}$ twisted bilayer graphene
It is known that Haldane model describes graphene in the limit when only the nearest neighbors couplings are present, i.e. at $t_{2}=0$. Graphene’s energy spectrum is gapless, yet is possesses ‘armchair‘ edge states at zero energy ([PhysRevB.84.195452, Wakabayashi_2010]). Numerically at $t_{inter}=0$, we observe edge states at zero energy at any kind of lattice termination. In addition, we observe states localized along one direction within the bulk at $E=\pm 1$ (see Fig. 20(a), 20(d)). The latter fact can be explained by writing graphene dispersion in momentum space as
$E=\pm\sqrt{\sin^{2}\left(\frac{\sqrt{3}k_{y}}{2}\right)+\left(\cos\left(\frac{\sqrt{3}k_{y}}{2}\right)+2\cos\left(\frac{k_{x}}{2}\right)\right)^{2}}$
(this equation can be obtained by writing the Hamiltonian for an isolated layer (2) at $t_{intra}=1$, $t_{2}=0$, $m=0$ and transforming it into momentum representation) and noticing that at $E=\pm 1$, its Fermi surface is hexagonal. In other words, at $E=\pm 1$, for the same value of $k_{x}$, there are many states with different $k_{y}$, which in turn means that states can become localized in $y$ direction.
Now let us explore the evolution of the energy spectrum with increasing interlayer coupling. We observe that at finite $t_{inter}$, the edge states get a finite energy width and in addition, get more localized (see Fig. 19). We also observe that a few of them turn into corner states. Finally we observe the emergence of new localized states at the center of the lattice (see Fig. 20(b) - 20(f)).
As the interlayer coupling increases even further, the edge states disappear, but instead a bulk gap opens (see Fig. 19). At $L_{x,y}=20,30$, we observe multiple in-gap states localized near edges. However, we observe corner states, as well as states localized at the center of the lattice at energies far away from the bulk gap (see Fig. 20(c), 20(f)).
For completeness, we also plot the dispersion as a function of $t_{2}$ for a fixed value of $t_{inter}$ (see Fig. 22). Specifically, at $t_{inter}=0$, we just observe that with decreasing $t_{2}$, the bulk gap gets closed. At $t_{inter}=1$, the bulk gap decreases with decreasing $t_{2}$, but does not close completely. Finally, the evolution becomes more non-trivial at large $t_{inter}$. Specifically at $t_{inter}=10$, we observe that as $t_{2}$ decreases from $1$ to 0, the bulk gap closes and then reopens again. In other words, we observe that the bulk gap at $30^{\circ}$ twisted bilayer graphene is not the same as the bulk gap in $30^{\circ}$ twisted Haldane model considered in the main text of the paper.
In summary, we have demonstrated that $30^{\circ}$ twisted bilayer graphene, similarly to the $30^{\circ}$ twisted bilayer Haldane model, hosts various states localized at the corners or at the center of the lattice. However, their energies are not attributed to the bulk gap, which opens at large interlayer coupling. This is another manifestation of the fact that the corner states are non-topological.
Figure 19: Energies of states in $30^{\circ}$ twisted bilayer graphene (i.e. the model from the Eqs. (1-3) with $t_{intra}=1$, $t_{2}=0$, $m=0$, $r_{0}=1$, $r_{max}=2$) plotted against $t_{inter}$. We consider lattice sizes $L_{x}=L_{y}=20,30,40$. The color represents the state’s fractal dimensions computed for $q=2$, as shown on the colorbar. We observe that at small non-zero $t_{inter}$, the edge states get finite energy width. At large $t_{inter}$, the edge states disappear, and a bulk energy gap opens.
((a))
((b))
((c))
((d))
((e))
((f))
Figure 20: Energies of states plotted against their fractal dimensions in $30^{\circ}$ twisted bilayer graphene (i.e. at $t_{2}=0$) for $m=0$, $L_{x,y}=30$, $t_{inter}=0$ (20(a)), $t_{inter}=1$ (20(b)), $t_{inter}=10$ (20(c)) and $L_{x,y}=40$, $t_{inter}=0$ (20(d)), $t_{inter}=1$ (20(e)), $t_{inter}=10$ (20(f)). One can see that at $t_{inter}=0$, the system hosts armchair edge states at zero energy, as well as localized bulk states due to band structure degeneracy. At $t_{inter}\neq 0$, the system hosts various localized states including states localized at the corners and at the center of the lattice respectively. We observe that the location of these states is not attributed to the bulk gap.
((a)) $t_{inter}=0.0$, $i=2757$, $E=0.0$
((b)) $t_{inter}=0.0$, $i=3441$, $E=1.0$
((c)) $t_{inter}=1.0$, $i=2024$, E=1.05E=-1.05
((d)) $t_{inter}=1.0$, $i=2727$, E=0.18E=-0.18
((e)) $t_{inter}=1.0$, $i=2801$, $E=0.11$
((f)) $t_{inter}=1.0$, $i=3114$, $E=0.24$
Figure 21: $|\psi_{i}|^{2}$ for a few of the states in $30^{\circ}$ twisted bilayer graphene (we choose $L_{x}=L_{y}=30$, and the other parameters are listed in the text of the Sec. A). We observe that at $t_{inter}=0$, the model hosts armchair edge states at $E=0$ (an example is shown on the Fig. 21(a)), as well as localized bulk states at $E=\pm 1$ (an example is shown on the Fig. 21(b)). On the other hand, at $t_{inter}=1$, the system still hosts edge states (21(d)), but in addition, it hosts states localized at the corners (21(e)), as well as at the center of the lattice (21(c), 21(f)).
((a)) $L_{x,y}=20$
((b)) $L_{x,y}=30$
((c)) $L_{x,y}=40$
Figure 22: Energies of states in the model determined by Eqs. (1 - 3) plotted against $t_{2}$. We assume $t_{intra}=1$, $m=0$, $r_{0}=1$, $r_{max}=2$. We also assume that the system has a shape of a square with $L_{x}=L_{y}$ and fix $L_{x,y}=20$ (22(a)), $L_{x,y}=30$ (22(b)) and $L_{x,y}=40$ (22(c)). We observe that at $t_{inter}=0$ the bulk gap closes at $t_{2}=0$, at $t_{inter}=1$, the bulk gap is decreasing at $t_{2}\to 0$ and at $t_{inter}=10$, the bulk gap closes and then reopens again.
Appendix B Calculation of the entanglement entropy
@@
Here we briefly describe the method we use to compute entanglement entropy, which was proposed in Refs. [9, 20] and used in a number of works later on, e.g. in [2]. We consider the model of twisted bilayer given by the Hamiltonian (1) and are interested in entanglement entropy of any subsystem from the Fig. 13. First, we compute eigenvectors $\psi$ of the Hamiltonian (1) and select only the ones corresponding to its filled states. Next, we compute ’correlation matrix’, i.e. sum of their outer products
$C_{ij}=\sum\limits_{f-filled\atop states}\psi_{f}(\vec{r}_{i})\psi_{f}^{\dagger}(\vec{r}_{j}).$ (11)
@@
After that we select the components whose coordinates $\vec{r}_{i},\vec{r}_{j}$ are inside the subsystem whose entanglement entropy we are interested in and thus obtain ’reduced correlation matrix’ $C_{A,ij}$. As it was shown in [9, 20], the entanglement entropy can be expressed in terms of its eigenvalues $\zeta_{m}$ as
@@
Finally, to substract the leading ’area law’ terms and extract the topological contribution, we compute the sum of entanglement entropies over several figures as shown on the Fig. 13.
Appendix C Kubo formula for anomalous Hall effect
@@
In this expression, we have transformed imaginary time into Matsubara frequencies representation, but left the coordinate representation. Here $w$ denote the Matsubara frequencies and $H_{ij}$ is the lattice Hamiltonian. The indices $i,j$ numerate the lattice sites. Electric field is introduced via Peierls substitution, which in turn may be approximated as
@@
$\vec{J}_{ij,\Omega}=i\sum\limits_{w}c^{\dagger}_{i,w+\Omega}H_{ij}(\vec{r}_{j}-\vec{r}_{i})c_{j,w}$ (15)
Note, that the current $\vec{J}_{ij,\Omega}$ at Matsubara frequency $\Omega$ is defined here as a vector for each bond connecting the lattice sites $i$, $j$.
Now we apply Kubo formula and write that the $x$ component of the expectation current has a form
$J_{x,ij,\Omega}=\langle J_{x,ij,\Omega}\sum\limits_{k.l}J_{y,kl,-\Omega}\rangle A_{y,kl,\Omega},$ (16)
where $A_{y,kl,\Omega}$ is a $y$ component of the vector potential at the bond connecting the lattice sites $k,l$ at Matsubara frequency $\Omega$. Now we assume that the electric field is constant in space, and the vector potential is written in the gauge $\vec{A}_{\Omega}=\vec{E}_{\Omega}/(i\Omega)$. The anomalous Hall conductivity is given by the antisymmetric part of the current-current correlator, namely
@@
Note that in our notations, the conductivity has just two indices $i,j$, which physically describe electric current at the bond connecting lattice sites $i,j$ as a response to constant electric field (if the electric field was not constant, the conductivity would have four components: two referring to the bond whose electric current we are interested and two referring to the bond whose electric field we look at).
We evaluate the current-current correlator in a conventional way by substituting fermionic Green’s functions and summing over Matsubara frequencies. The Green’s function can be written in terms of eigenstates $\psi_{n}(\vec{r}_{i})$ and energies $E_{n}$ (here $n$ numerates the states) as
$G_{ij,w}=\sum\limits_{n}\frac{\psi_{n}(\vec{r}_{i})\psi^{\dagger}_{n}(\vec{r}_{j})}{iw-E_{n}}.$ (18)
@@
Finally, let us look at the sum $\sum\limits_{ij}\sigma^{(Hall)}_{ij}$ over all lattice sites $i,j$. We can simplify the Eq. (II.3.3) by using explicitly the fact that $\psi_{n}(\vec{r}_{i})$ are eigenvectors, namely by applying an identity
@@
We remind that in these notations $\psi_{f,k}\equiv\psi_{f}(\vec{r}_{k})$ and $\psi_{e,k}\equiv\psi_{e}(\vec{r}_{k})$. Also in this equation, $x_{k}$ means $x$ coordinate of the lattice site $k$, and in the same way $x_{l}$ means $x$ coordinate of the lattice site $l$. We can obtain that the total sum has the form
$\sum\limits_{i,j}\sigma^{(Hall)}_{ij}=\frac{1}{2}\sum\limits_{f-filled\atop e-empty}\sum\limits_{i,k}$ (20)
@@
Here the summation is taken over filled states numerated by an index $f$ and empty states numerated by an index $e$, as well as lattice sites numerated by indices $k,l$. After applying the completeness relation
@@
Appendix D Anomalous Hall conductivity on an infinite crystalline lattice
The goal of this section is to show that the expression for Hall conductivity (II.3.3) from the main text in the case of an infinite crystalline lattice is indeed equivalent to its well-known expression in terms of Berry curvature. Here, for clarity we separate lattice indices $i,j.$... into pairs $(i,\alpha),(j,\beta),$..., which refer to unit cells and sublattice degrees of freedom respectively.
@@ Appendix D Anomalous Hall conductivity on an infinite crystalline lattice
$\psi_{n,i\alpha}=\frac{1}{\sqrt{N}}e^{i\vec{k}\vec{r}_{i}}\phi_{n,\alpha,\vec{k}}.$ (23)
Here the wavefunctions $\phi_{n,\alpha,\vec{k}}$ are eigenfunctions of the Schrodinger equation in momentum representation
$E_{n,\vec{k}}\phi_{n,\alpha,\vec{k}}=\sum\limits_{\beta}H_{\alpha\beta}(\vec{k})\phi_{n,\beta,\vec{k}}$ (24)
@@
$H_{\alpha\beta}(\vec{k})=\sum\limits_{j,\beta}e^{i\vec{k}(\vec{r}_{j}-\vec{r}_{i})}H_{i,j\atop\alpha,\beta}.$ (25)
@@
$H_{i,j\atop\alpha,\beta}=\frac{1}{N}\sum\limits_{\vec{k}}H_{\alpha\beta}(\vec{k})e^{-i\vec{k}(\vec{r}_{j}-\vec{r}_{i})}.$ (26)
@@
$\sum\limits_{j}\sigma^{(Hall)}_{i,j\atop\alpha,\beta}$ $=$ $\frac{1}{2}\sum\limits_{\gamma,\delta}\sum\limits_{f-filled\atop e-empty}\frac{1}{N}\sum\limits_{\vec{k}}\frac{1}{(E_{f,\vec{k}}-E_{e,\vec{k}})^{2}}$
$\times\left\{\phi^{\dagger}_{e,\alpha,\vec{k}}\frac{\partial H_{\alpha\beta}(\vec{k})}{\partial k_{x}}\phi_{f,\beta,\vec{k}}\cdot\phi^{\dagger}_{f,\gamma,\vec{k}}\frac{\partial H_{\gamma\delta}(\vec{k})}{\partial k_{y}}\phi_{e,\delta,\vec{k}}\right$.
$\qquad-\phi^{\dagger}_{f,\alpha,\vec{k}}\frac{\partial H_{\alpha\beta}(\vec{k})}{\partial k_{x}}\phi_{e,\beta,\vec{k}}\cdot\phi^{\dagger}_{e,\gamma,\vec{k}}\frac{\partial H_{\gamma\delta}(\vec{k})}{\partial k_{y}}\phi_{f,\delta,\vec{k}}$
@@
Notice that in the left side of this equation we have sum $\sum\limits_{j}\sigma_{ij}$ of conductivities through all bonds adjacent to a site $i$ - the quantity we discuss in the main text. In addition, to obtain the physical answer we have to sum over all sublattice degrees of freedom, i.e. over the indices $\alpha,\beta$. We also note that in the Eq. (D) we still assume that the lattice is discrete, but the momentum is continuum because the lattice is infinite. Hence it is straightforward to replace the momentum summation with an integration
$\frac{1}{N}\sum\limits_{\vec{k}}\to S\int\frac{d^{2}k}{(2\pi)^{2}},$ (28)
@@
$\sum\limits_{\alpha\beta}\phi^{\dagger}_{e,\alpha,\vec{k}}\frac{\partial H_{\alpha\beta}(\vec{k})}{\partial k_{x}}\phi_{f,\beta,\vec{k}}=\sum\limits_{\alpha}(E_{f,\vec{k}}-E_{e,\vec{k}})\phi^{\dagger}_{e,\alpha,\vec{k}}\frac{\partial\phi_{f,\alpha,\vec{k}}}{\partial k_{x}}$